# FedRSU: Federated Learning for Scene Flow Estimation on Roadside Units

### v1

@@ IV-B Dataset Properties
RSUs’ Setup. In total, the RSU-SF dataset comprises 17 Roadside Units, the specific configurations and parameters of each RSU are presented in Table I. It is evident that different RSUs exhibit remarkable diversity in various ways, including scenarios, modalities, LiDAR devices, and installation methods.
@@
Comparisons with existing datasets. Table II shows the comparisons among our constructed RSU-SF and other datasets. RSU-SF is a practical multi-modal FL dataset with point cloud and image modalities for scene flow estimation task. Unlike previous works that assume each client has labeled data, our RSU-SF represents more practical scenarios where data is generally unlabeled. Besides, previous works either emphasize on benchmarking generalized FL or personalized FL, while we simultaneously benchmark generalized and personalized FL. These properties also make RSU-SF a new and representative test-bed candidate for FL algorithms.
@@ V-A Implementation Details
Training. Each experiment is conducted based on PyTorch library on one Nvidia GeForce RTX 3090. We set the number of communication rounds as $50$. Within each round, each client trains the local model for $1$ epoch with a batch size of $32$. The optimizer used is ADAM with a learning rate of $4e^{-4}$. Unless specified, we adopt FlowStep3D [27] as the local scene flow estimation method. The hyper-parameters of loss terms are $\beta_{ch}=0.75$ and $\beta_{reg}=0.25$, respectively. For single-modal self-supervised learning, we discard the $s_{a}$ in loss function (1). Except for showing the effectiveness of our multi-modal method, we choose the single-modal version for benchmarking FedRSU to offer a clearer and more foundational perspective on this novel FL framework.
@@ V-C Benchmark: Generalized Federated Learning
Effects of the number of multi-modal clients in our multi-modal FL approach. We have shown the strong performance of our proposed multi-modal FL approach in TABLE III and IV. Further, we explore the effects of the ratio of multi-modal clients to provide deep understanding of the effectiveness of multi-modality. In TABLE VI, we randomly discard the image data of some clients, where the experiments are conducted on DAIR-V2X subset of RSU-SF to exclude the effects of other factors. From the table, we see that the performance generally degrades with the number of multi-modal clients decreases. This further verifies the effectiveness of our proposed multi-modal training.
@@ V-D Benchmark: Personalized Federated Learning
From Table VII, we see that (1) the gap between local learning and central learning with regard to mean value of epe3d is large, showing the need to augment individual RSU with outer information. (2) Most FL methods perform better than local learning, indicating that FL enables improving individual utility through collaboration. (3) Personalized FL methods that apply model partition (i.e., FedPer [54] and FedRep [55]) fail at this setting. This may due to that the patterns in convolutional neural networks (CNNs) for vision domain may not be applicable to our model architecture for lidar domain. In vision domain, there is clear evidence that shallow layers in a CNN capture more general patterns and thus may be more shareable in FL settings. However, this may not hold true in lidar domain, which may requires more explanation for future works and is not the focus in this paper. (4) Considering epe3d, pFedMe [56] not only achieves the highest utility, but also maintains the highest level of fairness. While pFedMe tends to has mediocre performance in previous literature in vision tasks [131, 57]. Again, this underlines the need for task-specific algorithms for this setting. Besides, this finding confirms the challenge of designing universal algorithms and calls for more future works to address such challenge.
@@ V-E Effects of Self-supervised Methods
V-F Visualization of Predicted Flow
@@ VI Discussion
VI-A Research Directions

### v2

@@ IV-B Dataset Properties
Clients’ setup. In total, the RSU-SF dataset comprises 17 roadside unit clients and 4 vehicle clients, the specific configurations and parameters of each RSU are presented in Table I. It is evident that different RSUs exhibit remarkable diversity in various ways, including scenarios, modalities, LiDAR devices, and installation methods.
@@
Comparisons with existing datasets. Table II shows the comparisons among our constructed RSU-SF and other datasets. RSU-SF is a practical multi-modal FL dataset with point cloud and image modalities for scene flow estimation task. Unlike previous works that assume each client has labeled data, our RSU-SF represents more practical scenarios where data is generally unlabeled. Besides, previous works either emphasize on benchmarking generalized FL or personalized FL, while we simultaneously benchmark generalized and personalized FL. These properties also make RSU-SF a new and representative test-bed candidate for FL algorithms. To the best of our knowledge, RSU-SF is the first real-world LiDAR-camera multi-modal dataset and benchmark for the federated learning community.
@@ V-A Implementation Details
Training. Each experiment is conducted based on PyTorch library on one Nvidia GeForce RTX 3090. We set the number of communication rounds as $50$. Within each round, each client trains the local model for $1$ epoch with a batch size of $32$. The optimizer used is ADAM with a learning rate of $4e^{-4}$. Unless specified, we adopt FlowStep3D [23] as the local scene flow estimation method. The hyper-parameters of loss terms are $\beta_{ch}=0.75$ and $\beta_{reg}=0.25$, respectively. For single-modal self-supervised learning, we discard the $s_{a}$ in loss function (1). Except for showing the effectiveness of our multi-modal method, we choose the single-modal version for benchmarking FedRSU to offer a clearer and more foundational perspective on this novel FL framework. Except for showing the effect of vehicle clients in our method, we use all 17 RSU clients in RSU-SF for experiments.
@@ V-C Benchmark: Generalized Federated Learning
Effects of the number of multi-modal clients in our multi-modal FL approach. We have shown the strong performance of our proposed multi-modal FL approach in TABLE III and IV. Further, we explore the effects of the ratio of multi-modal clients to provide deep understanding of the effectiveness of multi-modality. In TABLE VI, we randomly discard the image data of some clients, where the experiments are conducted on DAIR-V2X subset of RSU-SF to exclude the effects of other factors. From the table, we see that the performance generally degrades with the number of multi-modal clients decreases. This further verifies the effectiveness of our proposed multi-modal training. The improvement also suggests that camera sensors can substantially reduce the dependency on high-end LiDAR sensors, thereby in practical deployment, cutting costs while maintaining robust detection capabilities.
@@ V-D Benchmark: Personalized Federated Learning
From Table VII, we see that (1) the gap between local learning and central learning with regard to mean value of epe3d is large, showing the need to augment individual RSU with outer information. (2) Most FL methods perform better than local learning, indicating that FL enables improving individual utility through collaboration. (3) Personalized FL methods that apply model partition (i.e., FedPer [50] and FedRep [51]) fail at this setting. This may due to that the patterns in convolutional neural networks (CNNs) for vision domain may not be applicable to our model architecture for lidar domain. In vision domain, there is clear evidence that shallow layers in a CNN capture more general patterns and thus may be more shareable in FL settings. However, this may not hold true in the lidar domain, which may require more explanation for future works and is not the focus of this paper. (4) Considering epe3d, pFedMe [52] not only achieves the highest utility, but also maintains the highest level of fairness. While pFedMe tends to have mediocre performance in previous literature in vision tasks [128, 53]. Again, this underlines the need for task-specific algorithms for this setting. Besides, this finding confirms the challenge of designing universal algorithms and calls for more future works to address such challenge.
@@ V-E Effects of Self-supervised Methods
V-F Effects of Aggregating Vehicle Sensor Data
Although FedRSU focuses on federated learning on RSUs, technically, with suitable V2X communication technologies, vehicle sensors connected to the RSU could also participate in FedRSU training. Therefore, we explore the effects of incorporating vehicle sensor data into our FedRSU method. The experiments are conducted on the DAIR-V2X subset of the RSU-SF dataset to control for external factors. In Table IX, we gradually increase the number of vehicle clients involved in the FedRSU training (from V=0 to V=4) by randomly introducing vehicle client data from the RSU-SF dataset. Experimental results show that directly incorporating vehicle client data into the FedRSU method’s training negatively impacts the overall model performance. This is due to significant differences in data characteristics between RSU and vehicle sensors: First, the deployment of RSU and vehicle sensors differ substantially, as shown in Table I, including differences in height and orientation. Second, RSU sensors are statically deployed, while vehicles are mobile, leading to significant noise in the point cloud data collected by vehicle LiDAR sensors. Addressing these challenges and effectively integrating vehicle data into FedRSU training is a challenging problem, and we leave it to future research.
TABLE IX: Effects of the number of vehicle clients (V) in our FedRSU method, where FedAvg [42] is applied.
Vehicle clients V=0 V=1 V=2 V=3 V=4
Central Learning 14.59 - - - 18.09
FedRSU (FedAvg) [42] 17.28 18.75 20.69 20.09 19.80
V-G Visualization of Predicted Flow
@@ VI Discussion
VI-A Real-world Deployment in Large Scale
One concern may be the high cost of LiDAR sensors when deployed on a large scale. However, considering the following points, this is not a severe issue. First, using LiDAR sensors in the RSU setting is prevalent, including many datasets and algorithms developed based on these datasets, as discussed in Section II-A. The usage and promotion of RSU LiDARs in practical applications is a growing trend. Second, from a cost perspective, the price of long-range LiDAR sensors has been consistently decreasing. Commercially, RSU perception systems with LiDAR sensors are already available from companies like Leishen and Robosense. Third, our proposed multi-modality approach can reduce dependency on expensive LiDAR sensors. By including camera data, our method can enhance overall performance by 15%, which suggests that camera sensors can substantially reduce the dependency on high-end LiDAR sensors, thereby cutting costs while maintaining robust detection capabilities.
VI-B Research Directions
@@
Additionally, while the FedRSU system primarily considers RSU sensor data, vehicle sensors connected to RSUs can also be a substantial source of training data. Despite our dataset providing both RSU and vehicle data, incorporating vehicle data directly into FedRSU training does not effectively improve the model’s performance under the current method design. This lack of efficacy is mainly due to the differences in data characteristics between RSU and vehicle sensors. Bridging this gap and effectively combining these two types of data calls for additional methodological design.